OpenStudy (samsan9):

what is the equation of the parabola in vertex form with focus at (2 -4) and directrix y=-6? A.) (x+2)^2=1/4(y+4) B.) (x-2)^2=4(y+5) C.) (y+4)^2=-4(x-2) ' D.) (x-2)^2=-1/4(y+5)

OpenStudy (driftracer305):

k focus (2,-4) = (h,k)

OpenStudy (samsan9):

ok and then what?

OpenStudy (driftracer305):

(x – h)2 = 4p(y – k) .......... so (x-2)^2 = 4p(y+4)

OpenStudy (samsan9):

that isnt one of the options though?

OpenStudy (samsan9):

so A?

OpenStudy (driftracer305):

no no........ wait up....... something's wrong...... lemme double check

OpenStudy (samsan9):

ok ill wait

OpenStudy (driftracer305):

yep it's B........ it got it now.......m 5000% sure it is B

OpenStudy (driftracer305):

@samsan9 ........sorry i messed up earlier dude......

OpenStudy (driftracer305):

@samsan9 u there man?

OpenStudy (samsan9):

yeah dude sorry ok ill try it

OpenStudy (driftracer305):

no i got it......... it is B

OpenStudy (samsan9):

hell yeah thank you dude you get a fan and a medal :D

OpenStudy (driftracer305):

lol....... but dont u wanna knw how?? or u figured it out hmm??

OpenStudy (samsan9):

yes i would appreciate especially with a final coming up

OpenStudy (driftracer305):

k so focus is never (h,k) my first mistake........ nw a VERTEX is in between the focus and directrix..... here -5 between -6 and -4......... so p = absolute value of -6+5 = 1 and vertex = (2,-5) = (h,k)......... (x-2)^2 = 4(1)(y+5)